OpenStudy (anonymous):

Solve the equation. Show your work. 45 = 3b + 69

OpenStudy (anonymous):

b=6

OpenStudy (anonymous):

45 = 3b +69 45 - 69 = 3b -24 = 3b -24/3 = b -8 = b or b = -8

Parth (parthkohli):

-3b = 69 - 45 -3b = 25 3b = -25 b = -25/3

OpenStudy (anonymous):

@ParthKohli so close dude. LOL!

OpenStudy (anonymous):

@ParthKohli check it dude. :))

Parth (parthkohli):

Oh yeah. Wrong hhaa